OpenStudy (anonymous):

Evaluate log_5(1/625)

OpenStudy (lgbasallote):

1/625 = 5^(-4)

Parth (parthkohli):

This essentially says: \(\Large \color{MidnightBlue}{\Rightarrow 5^x = {1 \over 625} }\) Evaluate x, and that is the answer.

OpenStudy (lgbasallote):

\[\Large \log_5 (\frac{1}{625}) = \log_5 5^{-4}\]

OpenStudy (anonymous):

I think its just -4...

Parth (parthkohli):

It is. Correcto!